# ChatCFD: an End-to-End CFD Agent with Domain-specific Structured Thinking

### v1

@@ Abstract
Computational Fluid Dynamics (CFD) is critical for scientific and engineering advancements but is hindered by operational complexity and the need for extensive domain expertise. This paper presents ChatCFD, a large language model (LLM)-driven pipeline that automates computational fluid dynamics (CFD) workflows within the OpenFOAM framework, enabling users to configure and execute complex simulations from natural language prompts or published literature with minimal prior expertise. A key innovation is its ability to think structurally in database construction, configuration validation, and error reflection, which systematically integrates domain-specific CFD and OpenFOAM knowledge (e.g., solvers, turbulence models, file dependencies) with general LLMs, thereby enhancing accuracy and adaptability. More specifically, ChatCFD employs a four-stage workflow: (1) Knowledge Base Construction, creating a structured JSON database from OpenFOAM tutorials and manuals; (2) User Input Processing, guiding users via a multimodal interface (dialogue, documents, mesh files); (3) Case File Initialization, generating case files using a preprocessed knowledge base; and (4) Simulation Execution and Error Correction, running simulations and resolving errors with Retrieval-Augmented Generation (RAG). Powered by DeepSeek-R1 and DeepSeek-V3 models, a multi-agent architecture, and specialized OpenFOAM knowledge, ChatCFD features an interactive chat interface, hierarchical parameter extraction, and a robust error correction system addressing dimension mismatches, missing files, persistent errors, and general issues. Validation results demonstrate that ChatCFD autonomously reproduces published CFD literature results without human intervention, a challenging task involving complex, unseen configurations beyond basic CFD examples and general LLM capabilities. This achievement underscores the significant advancement enabled by its domain-specific structured thinking ability. In addition, validation experiments demonstrate error-free configurations in 30-40% of incompressible and compressible CFD cases and operational success in 60-80%, establishing a benchmark for automated CFD. By reducing expertise barriers, ChatCFD enhances CFD accessibility and provides methodological insights for AI-driven engineering simulations.
github: The code for ChatCFD is available at: https://github.com/EarlFan/ChatCFD.
Keywords large language model, multi-agent system, computational fluid dynamics, automated CFD
@@ 1 Introduction
Computational Fluid Dynamics (CFD) is a cornerstone technology in diverse scientific and engineering disciplines, including aerospace [28, 14, 12], energy systems [17, 2], urban environment [22, 19, 27], combustion [26, 10, 11], and biomedical applications [24, 9, 7]. It provides indispensable tools for simulating intricate fluid behaviors, thereby enabling design innovation and scientific discovery [3, 4]. However, the practical deployment of CFD faces significant challenges. Traditional CFD simulations require extensive domain-specific expertise [32] and frequently rely on costly commercial software packages. Even experienced engineers must invest considerable time in critical tasks, including solver selection, model development, mesh generation, boundary condition specification, and post-processing[20]. These complexities, compounded by the substantial computational resources needed, restrict the broader adoption of CFD, particularly for smaller enterprises and research groups, thus creating high entry barriers and potentially stifling innovation [4]. Consequently, there is a critical need for automated, user-friendly, and cost-effective CFD tools. To address this, we introduce ChatCFD, an automated agent system that leverages advanced artificial intelligence to streamline the CFD workflow.
Recent advances in Artificial Intelligence (AI) have revolutionized the automation landscape of complex scientific workflows. At the forefront of this transformation are Large Language Models (LLMs)—including GPT [1], Gemini [30], and DeepSeek [13]—alongside advanced multi-agent frameworks such as MetaGPT [15], and AutoGen [34]. These LLMs exhibit remarkable proficiency in natural language understanding, code generation, and complex reasoning tasks through chain-of-thought processes [33], enabling the development of autonomous LLM-based agents that can effectively perceive, reason, and act within their operational environment. Furthermore, the integration of Retrieval-Augmented Generation (RAG) [21] significantly enhances these agents. RAG is a technique where, before generating a response or performing an action, the LLM retrieves relevant information from an external, often domain-specific, knowledge base. This retrieved context is then provided to the LLM to inform its output. This process strengthens agents by grounding their responses in factual data, reducing hallucinations, and enabling them to perform more effectively in specialized domains like CFD.
This technological convergence has catalyzed the development of powerful LLM-based agents capable of automating the CFD workflow spectrum, from interpreting natural language specifications to configuring and executing simulations. This advancement represents a significant step toward making complex CFD simulations more accessible and efficient. OpenFOAM [18], in particular, stands out as the most popular open-source and free CFD software package, whose extensive capabilities, flexible architecture, and vibrant community support have made it a preferred choice for academic research, industrial applications, and consequently, a common framework for developing novel CFD automation agents. This accessibility has fostered innovation, leading to several pioneering initiatives exploring the integration of LLMs within the CFD domain. For example, MetaOpenFOAM [5] was introduced as an LLM-based multi-agent framework utilizing the MetaGPT [15] paradigm and RAG to automate CFD simulation tasks from natural language inputs, primarily demonstrated with OpenFOAM tutorial cases. Concurrently, OpenFOAMGPT [25] explored RAG-augmented LLM agents for OpenFOAM-centric CFD simulations, assessing model proficiency in tasks like zero-shot case setup and boundary condition adjustments. Foam-Agent [23] further advanced this area by proposing a multi-agent system with innovations such as hierarchical multi-index retrieval and dependency-aware file generation. Other research avenues have concentrated on fine-tuning LLMs with domain-specific datasets, like NL2FOAM [8], to enhance direct translation from natural language to executable CFD configurations, potentially using techniques like LoRA [16]. Despite these significant strides in automating CFD pipeline components and reducing technical barriers, existing systems often struggle with the end-to-end management of complex simulations without human intervention.
Existing CFD automation agents are often limited by their focus on rudimentary tasks, typically restricted to OpenFOAM tutorial-level examples, raising concerns about their generalization to complex, unseen cases. Three critical challenges remain unaddressed. First, describing CFD cases using brief natural language is often inadequate, as practical simulations require precise specifications beyond simplified verbal descriptions. If not given a detailed and professional task description, effective agents must interactively clarify user intent to define simulation requirements accurately. Second, handling intricate geometries poses significant difficulties, as mesh generation for real-world scenarios exceeds the capabilities of basic OpenFOAM utilities like “blockMesh” and demands specialized meshing tools. The inability to integrate user-defined or externally generated meshes limits these agents’ utility for diverse, practical CFD problems. Third, incorporating domain-specific knowledge into CFD agents remains challenging. Engineering-scale simulations require comprehensive initial and boundary condition specifications, necessitating domain-specific structural thinking and seamless integration of specialized domain expertise into the automation pipeline. These shortcomings highlight a critical gap in current CFD automation frameworks.
@@
To address the limitations outlined above, this paper introduces ChatCFD, an automated agent system for configuring and executing CFD simulations within the OpenFOAM framework, as illustrated in Fig. 1. ChatCFD enables users to define cases interactively via a chat interface or by uploading detailed technical files, such as well-documented research papers. Leveraging large language models (e.g., DeepSeek-R1, DeepSeek-V3), a sophisticated multi-agent architecture, and specialized OpenFOAM knowledge, ChatCFD autonomously manages the entire simulation workflow. It adeptly handles complex CFD setups through iterative trial-and-error, supporting comprehensive initial and boundary conditions, intricate physical models, and user-defined or externally generated mesh files. This architecture enables ChatCFD to address diverse, unseen scenarios and adapt various physical models, marking a significant advancement over prior agents.
This paper presents the detailed architecture of the ChatCFD system and validates its modular effectiveness through experimental analysis. The paper is organized as follows: Section 2 details the system architecture, Section 3 provides experimental results demonstrating the effectiveness of its components, and Section 4 summarizes the key findings. An appendix includes illustrative examples and prompts to clarify ChatCFD’s operational functionality.
2 Method
Figure 2: Architecture of the ChatCFD framework for automated CFD simulations, illustrating the four-stage workflow and agent structure. The stages are: (1) Knowledge Base Construction, creating a JSON database from OpenFOAM manuals and tutorials; (2) User Input Processing, enabling user interaction via natural language or document and mesh uploads; (3) Case File Initialization, generating OpenFOAM case files using the knowledge base; and (4) Simulation Execution and Error Reflection, running simulations, converting meshes with fluentMeshToFoam, and resolving errors (dimension mismatches, missing files, persistent errors, general issues) using RAG-based modules ReferenceRetriever and ContextRetriever. The agent structure integrates DeepSeek-R1 and DeepSeek-V3 for intelligent processing, with iterative error correction.
@@ 2.1 Stage 0: Knowledge Base Construction
Initially, preprocessed manual and tutorial data define the file dependency relationship, accessed during the RAG process to enhance the agent’s error correction ability. To ensure robust operation, strategic case filtering is implemented, removing OpenFOAM-specific headers, excluding cases with external dependencies or auxiliary folders to maintain independence, and filtering out cases with extensive non-uniform field definitions that reduce LLM response quality without contributing to configuration. Filtering criteria and implementation details are provided in Appendix A.
Subsequently, the example library is preprocessed through systematic extraction and organization. All case configuration files are parsed and converted into a structured JSON format, tagged with metadata such as solver specifications, turbulence model types, and flow regime characteristics. Analysis of solver and turbulence model distributions determines specific file requirements for each configuration. For instance, the simpleFoam solver requires configuration files including system/fvSolution, system/controlDict, system/fvSchemes, 0/p, 0/U, constant/transportProperties, and constant/turbulenceProperties. Similarly, the kOmegaSST turbulence model requires system/fvSolution, system/controlDict, system/fvSchemes, 0/p, 0/U, 0/k, 0/omega, 0/nut, constant/turbulenceProperties, and constant/transportProperties.
@@ 2.2 Stage 1: User Input Processing
  • Mesh Integration: The system assists users in uploading mesh files, currently supporting Fluent-format .msh files.
@@
2.3 Stage 2: Case File Initialization
@@
The CFD engineer layer, powered by the DeepSeek-R1 model, initializes OpenFOAM case files through a streamlined two-phase process, leveraging the preprocessed knowledge base.
In the first phase, the system generates a configuration list of required case files by analyzing the case description from Stage 1 to determine the appropriate solver and turbulence model. It queries the knowledge base’s OpenFOAM dataset to compile a tailored list of necessary files (e.g., system/fvSolution, 0/p) for the solver-turbulence model pair, enhancing initialization accuracy and minimizing missing file errors. Such a procedure, based on an explicit manner, substantially enhances the quality of case file initialization and reduces the incidence of missing file errors.
The second phase extracts detailed case setups from three sources: the configuration list, user prompt, and example library. The configuration list ensures precise case identification, while the prompt is segmented by section and filtered using the sentence-transformers/all-mpnet-base-v2 library with relevance thresholds. Hierarchical parameter extraction follows three steps:
  • Boundary Condition Validation: The system identifies boundary types from the prompt, verifies compatibility with OpenFOAM v2406, and enforces strict naming conventions to avoid errors. For example, in the airfoil case shown in Fig. 3, while the LLM may default to using a generic “inlet” type for the Inlet boundary based on common practice, the system ensures it correctly implements the specified “freeStream” boundary condition type to maintain fidelity with the original case setup.
  • Field Value Specification: Initial and boundary condition values are extracted for all field files (e.g., 0/p, 0/nut, 0/nuTilda) listed in the configuration.
  • Configuration File Generation: All required OpenFOAM files are generated, including field files (0/), physical properties (constant/, excluding polyMesh/), and control parameters (system/*).
Step 3 leverages DeepSeek-R1’s reasoning capabilities to orchestrate the simultaneous generation of all required case configuration files while ensuring consistency across physical models and boundary conditions. For numerical configurations not explicitly specified in the research article, such as discretization schemes and solution algorithms defined in system/fvSchemes and system/fvSolution files, ChatCFD employs DeepSeek-R1’s advanced reasoning capabilities to construct appropriate configurations based on its understanding of CFD best practices and the physical requirements of the case. The agent maintains coherence between interdependent parameters while adapting numerical schemes to match the flow physics and solution requirements indicated by the article’s methodology.
2.4 Stage 3: Simulation Execution and Error Reflection
Stage 3 executes the CFD simulation and analyzes runtime logs for automated error correction through three sequential steps:
  • Case Deployment: Case files from Stage 2 are deployed to the designated directory, and uploaded mesh files are converted to OpenFOAM format using fluentMeshToFoam, storing results in constant/polyMesh.
  • Temporal Configuration: Temporal parameters in system/controlDict are set, running 10 iterations/time steps for steady-state or transient simulations, with outputs at steps 0, 5, and 10. Transient cases enforce a maximum CFL number of 0.6, with time steps of $1\times 10^{-5}$ s for incompressible flows and $1\times 10^{-8}$ s for compressible flows.
  • Simulation Execution: The simulation runs using the specified OpenFOAM solver, with execution output logged to case_run.log for analysis.
ChatCFD automates error detection and correction in OpenFOAM case files by analyzing runtime logs using RAG and two retrieval modules:
The ReferenceRetriever module queries the preprocessed OpenFOAM knowledge database to select reference case files with matching or similar solvers and physical configurations. Typically, three reference files balance guidance and prompt complexity, ensuring contextually relevant corrections (e.g., distinguishing incompressible vs. compressible flows).
The ContextRetriever module compiles all current case configurations into a structured JSON format, providing the DeepSeek-V3 model with comprehensive context to address errors like missing files, especially in physically coupled problems (Section 3.3).
Upon detecting a runtime error, ChatCFD classifies it into four categories, detailed in Appendix B with examples:
  • Dimension Mismatches: Inconsistent physical dimensions in field or configuration files are identified by DeepSeek-V3 through log pattern analysis. Corrections involve adjusting dimensions to match reference files from ReferenceRetriever, ensuring consistency.
  • Missing Files: Missing case files, as reported in MetaOpenFOAM [5], are detected via log messages. DeepSeek-V3 identifies the missing file, ReferenceRetriever provides a template, and ContextRetriever supplies existing case details. DeepSeek-R1 generates the file, ensuring consistency in boundary patches, dimensions, and solver settings.
  • Persistent Errors: Errors recurring across correction cycles are tracked by DeepSeek-R1. The problematic file is identified, and DeepSeek-V3 rewrites it using reference files from ReferenceRetriever and the original file content.
  • General Errors: Errors like incorrect keywords or numerical schemes, comprising  50% of runtime issues, are addressed through a three-step process:
    • Error Localization: DeepSeek-R1 analyzes the log and case files (via ContextRetriever) to pinpoint the erroneous file, constraining the search to relevant files to reduce LLM errors.
    • Correction Proposal: DeepSeek-R1 generates step-by-step correction advice using the error log, erroneous file, and reference files from ReferenceRetriever, preserving initial and boundary conditions from Stage 1.
    • Correction Application: DeepSeek-V3 applies the correction, adhering to the advice and maintaining dimensional consistency.
2.5 LLM Configurations
ChatCFD employs a dual-model approach, integrating DeepSeek-R1 and DeepSeek-V3 to handle the initiation and correction of case files in the workflow. The application of these two LLMs follows distinct principles. DeepSeek-R1 takes the lead in Stage 2, where it extracts comprehensive case setups from textual descriptions and configures the initial case files. Its capacity for in-depth text comprehension and error correction is also crucial in Stage 3, where it analyzes and rectifies general errors and generates any missing files. DeepSeek-V3’s contributions are concentrated in Stage 3, focusing on the correction of case files. It excels at managing highly patterned tasks, such as identifying dimensional and file-missing errors. Furthermore, DeepSeek-V3 handles specific corrections, such as adjusting field file dimensions, rebuilding case files to correct persistent errors, and amending files according to precise step-by-step instructions.
This synergistic use of DeepSeek-R1 and DeepSeek-V3 significantly enhances ChatCFD’s overall performance, particularly in correcting case files. While DeepSeek-R1 excels in Stage 3 with its strong reasoning and planning for long-text comprehension and comprehensive file configuration, its tendency to hallucinate or generate inaccuracies became a hurdle during development. These hallucinations led to problems such as misidentifying dimensional errors and inserting unwanted markdown formatting into corrected files, issues that extensive prompt engineering alone could not resolve.
To address these challenges, ChatCFD adopted a strategic hybrid method. In the case file correction module, DeepSeek-V3 now manages the majority of error corrections, which has markedly increased the success rate of these corrections. DeepSeek-R1 retains responsibility for resolving file-missing errors, as this task requires a holistic analysis of all configuration files for consistency—a strength of DeepSeek-R1’s content processing. A similar division applies to error type identification: DeepSeek-V3 handles pattern-based errors like dimension and file-missing issues, while DeepSeek-R1 tackles complex errors, such as repetitive and common problems, which necessitate broader content analysis.
For DeepSeek-V3, the thought prompting by Wu et al. [35] in 2024 is used to equip DeepSeek-V3 with some thinking abilities. An auxiliary function then helps to extract the final answer from DeepSeek-V3’s response. The thought prompting text instructs the model to respond comprehensively and in detail, and to first write down its thought process after the phrase "Here is my thought process:" before providing its actual response after "Here is my response:". An example of thought prompting is given in Appendix C.
Furthermore, ChatCFD maintains a log of all Question-Answer (QA) histories with the LLMs for each CFD case. These QA logs are valuable for analyzing the LLM’s behavior as it configures CFD cases within the OpenFOAM framework.
The experimental validation detailed in Section 3 was conducted on the VolcEngine platform [31] from March 1 to March 12, 2025. The pricing structure for the models on this platform, used for evaluating token usage and cost, is as follows: for DeepSeek-R1, the cost is \$0.00055 (0.004 RMB) per 1,000 input tokens and \$0.0022 (0.016 RMB) per 1,000 output tokens; for DeepSeek-V3, the cost is \$0.00021 (0.0015 RMB) per 1,000 input tokens and \$0.00082 (0.006 RMB) per 1,000 output tokens. All token usage and cost calculations mentioned in Section 3 are based on this pricing structure from the VolcEngine platform.

### v2

# ChatCFD: An LLM-Driven Agent for End-to-End CFD Automation with Domain-Specific Structured Reasoning

@@ Abstract
Computational Fluid Dynamics (CFD) is essential for advancing scientific and engineering fields, but is hindered by operational complexity, high expertise requirements, and limited accessibility. This paper introduces ChatCFD, an automated agent system for OpenFOAM simulations that processes multi-modal inputs (e.g., research papers, meshes) via an interactive interface, leveraging DeepSeek-R1 and DeepSeek-V3 large language models (LLMs), a multi-agent architecture, and OpenFOAM knowledge. Its four-stage pipeline, Knowledge Base Construction, User Input Processing, Case File Generation, and Execution and Error Reflection, enables iterative trial-reflection-refinement for intricate setups, supporting diverse physical models and external meshes. Validation on 205 benchmark tutorial cases, 110 perturbed variants, and 2 literature-derived cases demonstrates ChatCFD’s 82.1% operational success rate on basic cases, significantly outperforming MetaOpenFOAM (6.2%) and Foam-Agent (42.3%), and 60–80% on literature-derived complex cases, with turbulence model studies showing a 40% success rate for common models vs. 10% for rare ones like RNGk-$\epsilon$. Physics coupling analyses reveal higher resource demands for multi-physics-coupled cases, while LLM bias toward simpler setups introduces persistent errors, such as dimensional inconsistency. Ablation studies highlight the efficacy of RAG-based modules and reflection mechanisms. By automating hypothesis testing and parameter exploration, ChatCFD accelerates scientific discovery in fluid mechanics and engineering, addressing LLM limitations through structured design and demonstrating strong potential as a modular component in MCP-based agent networks for collaborative multi-agent systems, paving the way for scalable AI-driven CFD innovation. The code for ChatCFD is available at: https://github.com/ConMoo/ChatCFD.
Keywords large language model, multi-agent system, computational fluid dynamics, OpenFOAM, automated CFD
@@ 1 Introduction
Computational Fluid Dynamics (CFD) is a cornerstone technology in diverse scientific and engineering disciplines, including aerospace [1, 2], energy systems [3, 4], urban environment [5, 6, 7], combustion [8, 9], and biomedical applications [10, 11]. It provides indispensable tools for simulating intricate fluid behaviors, thereby enabling design innovation and scientific discovery [12, 13]. However, CFD implementation faces substantial barriers, requiring deep domain expertise [14] and often relying on expensive commercial software. Even experts invest significant time in tasks such as solver selection, model setup, mesh generation, boundary condition definition, and post-processing [15]. These demands, combined with high computational costs, limit CFD accessibility for smaller organizations and hinder broader innovation [13]. Thus, there is an urgent need for automated, intuitive, and affordable CFD solutions. To meet this need, we present ChatCFD, an AI-driven agent system that streamlines CFD workflows, facilitating automated scientific discovery in fluid mechanics and engineering by enabling rapid iteration and exploration of complex flow phenomena without extensive manual intervention.
Recent progress in Artificial Intelligence (AI) has transformed the automation of sophisticated scientific processes. Large Language Models (LLMs), such as GPT [16], Gemini [17], and DeepSeek [18], alongside multi-agent frameworks like MetaGPT [19] and AutoGen [20], excel in natural language processing, code generation, and reasoning via chain-of-thought techniques [21]. Frameworks like ReAct [22] synergize reasoning and acting in LLMs, while Reflexion [23] uses verbal reinforcement learning to improve agent performance through self-reflection. Toolformer [24] teaches LLMs to self-learn tool usage. Retrieval-Augmented Generation (RAG) [25] further bolsters these agents by incorporating domain-specific knowledge, reducing hallucinations, and improving domain adaptation for tasks like CFD automation.
This synergy has enabled LLM-based agents to automate CFD pipelines, from interpreting user inputs to running simulations. OpenFOAM [26], a leading open-source CFD platform, supports this trend due to its flexibility and community backing. Notable efforts include MetaOpenFOAM [27], which automates simulations from natural language using RAG and MetaGPT, though limited to tutorial-level cases and struggling with complex geometries; OpenFOAMGPT [28], evaluating RAG-augmented LLMs for zero-shot setup and boundary adjustments, but facing difficulties in turbulence model modifications; Foam-Agent [29], introducing hierarchical retrieval for dependency-aware generation, yet constrained in handling unseen configurations; and NL2FOAM [30], fine-tuning LLMs for natural language-to-CFD translation using LoRA [31], but reliant on domain-specific datasets and less effective for diverse physical models. Despite these advances, existing systems often fall short in end-to-end automation for complex cases without human oversight, particularly due to a fundamental challenge: LLMs lack sufficient training on niche scientific tasks like OpenFOAM setup, necessitating expert-designed architectures to bridge this gap through structured reasoning, reflection, and knowledge integration.
Existing CFD automation agents are often limited by their focus on rudimentary tasks, typically restricted to OpenFOAM tutorial-level examples, raising concerns about their generalization to complex, unseen cases. This challenge primarily stems from the scarcity of domain-specific scientific coding corpora, such as OpenFOAM setups, requiring sophisticated multi-agent system designs to effectively incorporate physical knowledge into agent architectures, thereby enhancing LLMs’ capabilities in specialized scientific tool invocation. Recent representative work in this area includes SciToolAgent [32], which develops a knowledge graph-driven agent for integrating hundreds of scientific tools across biology, chemistry, and materials science, using graph-based RAG and safety-checking to automate workflows like protein engineering and chemical synthesis. While not directly focused on CFD, its approach to tool integration and domain adaptation offers valuable insights for fluid automation agents in handling coupled physics and complex simulations.
Three critical challenges in CFD automation agents remain unaddressed. First, incorporating domain-specific knowledge into CFD agents remains challenging. Engineering-scale simulations require comprehensive initial and boundary condition specifications, necessitating domain-specific structural thinking and seamless integration of specialized domain expertise into the automation pipeline. Second, designing effective agent frameworks, especially for future multi-MCP agent collaborations, is essential to maximize complementary capabilities. Typical issues include whether we need to contain mesh generation in the agent scope, where real-world geometries exceed basic OpenFOAM tools like blockMesh and require specialized integration of external meshes. Third, enabling large-scale testing and evolution of fluid agents is crucial. Current reliance on simplistic natural language descriptions often leads to ambiguous case setups and hinders scalability. In contrast, CFD literature, comprising millions of papers with detailed descriptions and standard results, remains underutilized, impeding agent evolution. These shortcomings highlight a critical gap in current CFD automation frameworks.
@@
To overcome these challenges, we introduce ChatCFD, an automated agent system for OpenFOAM simulations, as shown in Figure 1. ChatCFD processes multi-modal inputs (e.g., research papers, meshes) through an interactive interface, utilizing LLMs (DeepSeek-R1, DeepSeek-V3), multi-agent architecture, and OpenFOAM knowledge to manage the full workflow. It supports intricate setups via iterative refinement, handling diverse physical models and external meshes, surpassing prior agents in adaptability to unseen cases.
This paper details ChatCFD’s architecture and validates its components experimentally. The structure is as follows: Section 2 describes the pipeline design, Section 3 presents results, and Section 4 summarizes findings. Appendices provide examples and prompts for operational clarity.
2 ChatCFD Pipeline Design
Figure 2: Architecture of the ChatCFD framework for automated CFD simulations, illustrating the four-stage workflow and agent structure. The stages are: (1) Knowledge Base Construction, creating a JSON database from OpenFOAM manuals and tutorials; (2) User Input Processing, enabling user interaction via natural language or document and mesh uploads; (3) Case File Generation, generating OpenFOAM case files using the knowledge base; and (4) Execution and Error Reflection, running simulations, converting meshes with fluentMeshToFoam, and resolving errors (dimension mismatches, missing files, persistent errors, general issues) using RAG-based modules ReferenceRetriever and ContextRetriever. The agent structure integrates DeepSeek-R1 and DeepSeek-V3 for intelligent processing, with iterative error correction.
@@ 2.1 Stage 0: Knowledge Base Construction
Initially, preprocessed manual and tutorial data are used to define file dependency relationships, which are accessed during the RAG process to enhance the agent’s error correction capabilities. To ensure robust operation, strategic case filtering is implemented by removing OpenFOAM-specific headers and excluding cases with external dependencies or auxiliary folders. This maintains case independence, reducing the risk of degraded LLM response quality due to irrelevant configuration elements. For cases where non-uniform field definitions occupy significant file content and may exceed the dialogue context limits of large models, placeholder tokens such as __CELL_COUNT__,__FACE_COUNT__,__SCALAR_DATA_PLACEHOLDER__ are used to replace non-uniform field data. This approach enables the LLM to focus on critical settings, such as boundary and initial conditions, while ignoring non-uniform field configurations. As a result, it mitigates the reduction in the number of usable OpenFOAM tutorial cases caused by complex non-uniform field data. Filtering criteria and implementation details are provided in Appendix A. Subsequently, the example library is preprocessed through systematic extraction and organization. All case configuration files are parsed and converted into a structured JSON format, tagged with metadata such as solver specifications, turbulence model types, and flow regime characteristics. Analysis of solver and turbulence model distributions determines specific file requirements for each configuration. For instance, the simpleFoam solver requires configuration files including system/fvSolution, system/controlDict, system/fvSchemes, 0/p, 0/U, constant/transportProperties, and constant/turbulenceProperties. Similarly, the k-$\omega$ SST turbulence model requires system/fvSolution, system/controlDict, system/fvSchemes, 0/p, 0/U, 0/k, 0/omega, 0/nut, constant/turbulenceProperties, and constant/transportProperties.
@@ 2.2 Stage 1: User Input Processing
  • Mesh Integration: The system assists users in uploading mesh files, currently supporting Fluent format .msh files. For OpenFOAM default meshes, users must specify the path to the corresponding polyMesh directory.
@@
2.3 Stage 2: Case File Generation
@@
The CFD engineer layer, powered by the DeepSeek-R1 and DeepSeek-V3 models, initializes OpenFOAM case files through a streamlined three-phase process that leverages the preprocessed knowledge base.
2.3.1 Phase 1: File Configuration Synthesis
The system analyzes the case description from Stage 1 (User Input Processing) to identify suitable solvers and models, thereby generating a list of required case file configurations. This involves explicitly matching dependencies for solvers, models, and related settings to facilitate accurate file configuration. The workflow proceeds as follows:
  • Relevant cases are retrieved based on the solver and model, yielding a corresponding file configuration list from which non-essential files (e.g., those for visualization or mesh generation) are removed. These refined cases are then input to the DeepSeek-V3 model to produce an optimized file configuration list.
  • In the absence of direct matches, cases are aligned by solver type. For additional models, such as turbulence models, necessary files are inferred from analogous cases, synthesized into a new configuration list, verified for solver compatibility, and finalized via DeepSeek-V3.
This approach integrates solver and physical model structures with case synthesis, enhancing initialization precision, minimizing file omissions, and substantially improving overall case file quality.initialization.
2.3.2 Phase 2: Case Setup Extraction
Detailed case configurations are derived from three primary sources: the file configuration list, user prompts, and the case library. The configuration list prevents omissions or redundancies in the case structure, while user prompts enable filtering of segmented case description chunks using the sentence-transformer/all-mpnet-base-v2 library and a similarity threshold. The DeepSeek-R1 model implements a hierarchical three-step extraction:
  • Mesh Boundary Condition Extraction: Boundary names and types are extracted via regular expressions and tools like pyFoam, forming a persistent dictionary for subsequent physical field generation and error handling.
  • Boundary Condition Validation: Boundary conditions are identified from filtered chunks, checked for compatibility with OpenFOAM-v2406, and subjected to rigorous naming and format validation.
  • Physical Field File Value Setup: Field files (e.g., 0/p, 0/nut, 0/nutTilda) are identified from the configuration list. Using prior boundary data, OpenFOAM-compliant templates are created, encompassing internal and boundary field settings differentiated by boundary names and associated keywords.
2.3.3 Phase 3: Configuration Validation and Refinement
For unspecified parameters, such as discretization schemes and solver algorithms in system/fvSolution and system/fvSchemes, the DeepSeek-R1 model applies advanced reasoning grounded in CFD best practices and case-specific physics to formulate coherent configurations. This maintains parameter interdependencies and ensures alignment with the outlined flow dynamics and solver demands.
A comprehensive validation and correction mechanism further refines the generated files. The system scrutinizes structures in the system and constant directories, confirming proper dependencies and completeness through cross-references with library exemplars. Particular emphasis is placed on physical field dimensions to mitigate LLM limitations in interpreting such constraints. For example, variables like p and alphat—which differ dimensionally between compressible and incompressible flows—are calibrated according to case type; ambiguous cases retain initial values. This process curtails errors like dimensional inconsistencies, bolsters initial file robustness, and diminishes iterative corrections in Stage 3.
2.4 Stage 3: Error Correction and Reflection
Stage 3 of the ChatCFD pipeline employs a modular architecture for automated error correction and iterative refinement in OpenFOAM case configurations. The following modules facilitate robust error handling:
  • ReferenceRetriever: This module retrieves reference files from the preprocessed OpenFOAM knowledge base, matching solver and model specifications from benchmark tutorial cases. If no exact match is found, it prioritizes solver-compatible files. To balance guidance and prompt complexity, two reference files are selected to aid error detection and correction, leveraging the LLM’s few-shot learning to enhance ChatCFD’s accuracy in error localization and resolution.
  • ContextRetriever: This module compiles current case configurations into a structured JSON format, providing detailed file content and modification trajectories over a specified period, tailored to other modules’ needs. It supports targeted reflection and correction, particularly for errors involving physical coupling (e.g., pressure-density interactions).
  • Error Localization Module: Comprising two components: (1) DeepSeek-V3 rapidly identifies suspicious files based on error messages, reducing the analysis burden; (2) DeepSeek-R1 performs detailed reasoning to pinpoint erroneous files by comparing their content against the case library, leveraging its advanced inference capabilities for precise localization.
  • Error Correction Module: This module addresses file interdependencies in OpenFOAM through a three-step process: (1) DeepSeek-V3 identifies files requiring coordinated modifications to resolve coupling-related errors; (2) DeepSeek-R1 proposes actionable corrections by integrating error details, related file content, simulation requirements, and benchmark tutorial cases; (3) DeepSeek-V3 applies these corrections, ensuring compliance with OpenFOAM’s file format standards.
  • Reflection Module: Activated during persistent errors, this module collects error messages and file modification histories, formulating reflections in the format: ”For situation A, I considered B but overlooked C; next, I will apply D to resolve the issue.” These reflections are stored as reflection blocks within a <reflexion> tag in the reflection history, enhancing iterative error correction by integrating with localization and correction modules.
Errors detected during simulation are categorized into two types:
  • General Errors: These, comprising over 70% of runtime errors, include incorrect keywords, formatting issues, or floating-point errors. Correction involves a streamlined workflow: suspicious file detection (DeepSeek-V3), error confirmation (DeepSeek-R1), retrieval of related tutorial files (ReferenceRetriever), proposal of corrections, and file modification. For missing file errors, DeepSeek-V3 directly generates the required file using reference cases, followed by dimensional validation to ensure quality.
  • Persistent Errors: ChatCFD leverages short-term memory (recent error messages and file modification histories) and long-term memory (reflection histories) to address recurring issues. Reflection histories, stored as structured insights, enhance the system’s ability to adapt and resolve complex errors iteratively.
2.5 LLM Architecture and Configuration
ChatCFD employs a dual-model architecture, integrating the DeepSeek-R1 and DeepSeek-V3 large language models (LLMs) to execute a comprehensive CFD workflow for case generation and error correction. Each model is strategically deployed based on its strengths to optimize performance across Stages 2 and 3.
The DeepSeek-R1 model, with its superior text comprehension and reasoning capabilities, is utilized for complex, high-stakes tasks. In Stage 2, it extracts critical simulation parameters (e.g., boundary conditions, initial conditions) from case descriptions and generates complete initial case files. In Stage 3, it pinpoints error-causing files, proposes detailed correction strategies, and conducts reflection on persistent errors. However, DeepSeek-R1’s propensity for hallucinations—such as inserting unnecessary Markdown formatting, which can disrupt error correction—poses challenges. Despite mitigation efforts using prompt engineering and pydantic for structured outputs, these issues persist. Additionally, its inference time increases significantly with longer prompts, making it less efficient for routine tasks.
The DeepSeek-V3 model excels in instruction following and rapid response, making it ideal for structured, low-complexity tasks. In Stage 2, it performs quick file structure validation, while in Stage 3, it identifies suspicious or missing files and generates corrected file content based on provided instructions. To address DeepSeek-V3’s limited reasoning capacity, tasks are simplified into structured formats (e.g., providing configuration file lists to narrow analysis scope), and the model is prompted to justify its actions (e.g., explaining why a file is suspicious). This enhances its effectiveness while maintaining computational efficiency. This dual-model approach leverages DeepSeek-R1’s reasoning for complex CFD tasks and DeepSeek-V3’s efficiency for structured operations, ensuring robust performance across diverse case types (e.g., benchmark tutorial cases, literature-derived cases) while minimizing computational costs.
ChatCFD maintains a comprehensive log for each CFD case, capturing all question-answer interactions and actions taken during reflection-based error correction. These logs are invaluable for analyzing the behavior of the DeepSeek-R1 and DeepSeek-V3 models within the OpenFOAM framework, particularly in identifying limitations in handling complex CFD configurations. The recorded reflection iterations, which document corrective actions for persistent errors, provide critical insights into scenarios where ChatCFD’s performance is suboptimal, guiding targeted improvements in the pipeline. Additionally, a concise summary of reflections for recurring errors is maintained, offering a deeper understanding of ChatCFD’s behavior in complex scenarios. These logs have the potential to form a novel database, serving as a valuable resource for enhancing LLM-based agents in addressing high-complexity CFD problems within the MCP framework.
The experimental validations described in Section 3 were conducted on the VolcEngine platform [33], which provides the pricing structure for token usage and computational costs. For DeepSeek-R1, input tokens cost \$0.00055 per 1,000 (0.004 RMB), and output tokens cost \$0.0022 per 1,000 (0.016 RMB). For DeepSeek-V3, input tokens cost \$0.00021 per 1,000 (0.0015 RMB), and output tokens cost \$0.00082 per 1,000 (0.006 RMB). All token consumption and cost metrics reported in Section 3 are based on this pricing, enabling precise evaluation of ChatCFD’s computational efficiency across benchmark tutorial cases, perturbed variant cases, and literature-derived cases.